# Copula-like inference for discrete bivariate distributions with rectangular supports

### v2

@@ 4.3 Monte Carlo experiments
The asymptotic results stated in Corollary 4.6 and Proposition 4.11 do not provide any information on the finite-sample behavior of the three method-of-moments estimators in (4.14) and the maximum pseudo-likelihood estimator in (4.17). To compare these four estimators, we carried out Monte Carlo experiments in the situation when the parametric family $\mathcal{J}$ in (4.11) is constructed from a one-parameter family of classical copulas as in (4.12). For $(r,s)\in\{(3,3),(3,10),(10,10)\}$, the bias and the mean squared error (MSE) of the three method-of-moments estimators and the maximum pseudo-likelihood estimator were estimated from 1000 random samples of size $n\in\{100,500,1000\}$ generated from a p.m.f. with copula p.m.f. of the form (4.12) with $C_{\theta}$ either the Clayton or the Gumbel–Hougaard copula with a Kendall’s tau in $\{0.33,0.66\}$, or the Frank copula with a Kendall’s tau in $\{-0.5,0,0.5\}$ (see, e.g., Hofert et al., 2018, Chapter 2 for the definitions of these copula families and the definition of Kendall’s tau). For each of the above seven copula p.m.f.s, three marginal scenarios were considered:
@@
    the uniform distribution on $I_{r}$ as first margin and the uniform distribution on $I_{s}$ as second margin,
@@
    $(1,\dots,r)/(r(r+1)/2)$ as first margin and $(1,\dots,s)/(s(s+1)/2)$ as second margin,
@@
    the binomial distribution with parameters $r-1$ and $1/2$ as first margin and the binomial distribution with parameters $s-1$ and $1/2$ as second margin.
@@ 5.3 Chi-square tests based on a semi-parametric bootstrap
    Let $p^{[1,n]}$ and $p^{[2,n]}$ denote the margins of $p^{[n]}$ in (4.2), and, using (2.5) with $a=p^{[1,n]}$ and $b=p^{[2,n]}$, form $p^{[n,\hat{\theta}^{[n]}]}=\mathcal{I}_{p^{[1,n]},p^{[2,n]}}(u^{[\hat{\theta}^{[n]}]})$, which is a consistent semi-parametric estimate of the p.m.f. of $(X,Y)$ under $H_{0}$ in (5.1).
@@ 6 Data example
We briefly illustrate the proposed methodology on a bivariate data set initially analyzed in Goodman (1979). The underlying discrete random variables are $X$, the occupational status of a British male subject, and $Y$ the occupational status of the subject’s father. The occupational status can take $r=s=8$ different ordered values encoded as the first 8 integers. Goodman (1979) does not explain why the underlying categories are considered to be ordered and refers the reader to Miller (1960), among others. We note that there seems to be no reasons to consider that certain values in $I_{r,s}$ cannot be realizations of $(X,Y)$. In other words, it seems meaningful to assume that the unknown p.m.f. $p$ of $(X,Y)$ has support $I_{r,s}$ and thus that the decomposition in (3.2) holds. The random vector $(X,Y)$ was observed on $n=3498$ subjects. The resulting bivariate contingency table can be obtained in R by typing data(occupationalStatus) and is reproduced below for convenience:

### v4

@@ 4.3 Monte Carlo experiments
The asymptotic results stated in Corollary 4.6 and Proposition 4.11 do not provide any information on the finite-sample behavior of the three method-of-moments estimators in (4.14) and the maximum pseudo-likelihood estimator in (4.17). To compare these four estimators, we carried out Monte Carlo experiments in the situation when the parametric family $\mathcal{J}$ in (4.11) is constructed from a one-parameter family of classical copulas as in (4.12). For $(r,s)\in\{(3,3),(3,10),(10,10)\}$, the bias and the mean squared error (MSE) of the three method-of-moments estimators and the maximum pseudo-likelihood estimator were estimated from 1000 random samples of size $n\in\{100,500,1000\}$ generated from a p.m.f. with copula p.m.f. of the form (4.12) with $C_{\theta}$ either the Clayton or the Gumbel–Hougaard copula with a Kendall’s tau in $\{0.33,0.66\}$, or the Frank copula with a Kendall’s tau in $\{-0.5,0,0.5\}$ (see, e.g., Hofert et al., 2018, Chapter 2 for the definitions of these copula families and the definition of Kendall’s tau). For each of the above seven copula p.m.f.s, three marginal p.m.f. scenarios were considered:
@@
    $(1/r,\dots,1/r)$ (resp. $(1/s,\dots,1/s)$) as values of the first (resp. second) marginal p.m.f.,
@@
    $(1,\dots,r)/(r(r+1)/2)$ (resp. $(1,\dots,s)/(s(s+1)/2)$) as values of the first (resp. second) marginal p.m.f.,
@@
    the values of the p.m.f. of the binomial distribution with parameters $r-1$ and $1/2$ (resp. $s-1$ and $1/2$) as values of the first (resp. second) marginal p.m.f.
@@ 5.3 Chi-square tests based on a semi-parametric bootstrap
    Let $p^{[n,1]}$ and $p^{[n,2]}$ denote the margins of $p^{[n]}$ in (4.2), and, using (2.5) with $a=p^{[n,1]}$ and $b=p^{[n,2]}$, form $p^{[n,\hat{\theta}^{[n]}]}=\mathcal{I}_{p^{[n,1]},p^{[n,2]}}(u^{[\hat{\theta}^{[n]}]})$, which is a consistent semi-parametric estimate of the p.m.f. of $(X,Y)$ under $H_{0}$ in (5.1).
@@ 6 Data example
We briefly illustrate the proposed methodology on a bivariate data set initially analyzed in Goodman (1979). The underlying discrete random variables are $X$, the occupational status of a British male subject, and $Y$ the occupational status of the subject’s father. The occupational status can take $r=s=8$ different ordered values encoded as the first 8 integers. Goodman (1979) does not explain why the underlying possible values are considered to be ordered and refers the reader to Miller (1960), among others. We note that there seems to be no reasons to consider that certain values in $I_{r,s}$ cannot be realizations of $(X,Y)$. In other words, it seems meaningful to assume that the unknown p.m.f. $p$ of $(X,Y)$ has support $I_{r,s}$ and thus that the decomposition in (3.2) holds. The random vector $(X,Y)$ was observed on $n=3498$ subjects. The resulting bivariate contingency table can be obtained in R by typing data(occupationalStatus) and is reproduced below for convenience: